# Efficient Diffusion Models: A Survey

### v1

@@ 3.3 Efficient Sampling
Efficient sampling in diffusion models addresses the computational challenges of the iterative denoising process while maintaining the quality of the generated samples through three main approaches. As illustrated in Figure 13, these encompass efficient SDE and ODE solvers, sampling scheduling strategies including parallel sampling and timestep optimization, and truncated sampling methods leveraging early exit and retrieval-based techniques.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=18em,font=,, where level=2text width=15em,font=,, where level=3text width=24em,font=,, [ Efficient Sampling, ver [ Efficient Solver, fill=blue!10 [ SDE Solver, fill=blue!10 [ DiffFlow (Zhang & Chen, 2021), DiNoF (Zand et al., 2023),  Jolicoeur-Martineau et al. (2021),
GMS (Guo et al., 2024), BFM-Solver2 (Xue et al., 2024a), SA-Solver (Xue et al., 2024b), leaf, text width=45em ] ] [ ODE Solver, fill=blue!10 [ DDIM (Song et al., 2020a), DPM-solver (Lu et al., 2022), DEIS (Zhang & Chen, 2022),
@@
TAA (Walker & Ni, 2011), leaf, text width=45em ] ] [ Timestep Scheduling, fill=blue!10 [ FastDPM (Kong & Ping, 2021),  Watson et al. (2021), AYS (Sabour et al., 2024), leaf, text width=45em ] ] ] [ Truncated Sampling, fill=blue!10 [ Early Exit, fill=blue!10 [ CATs (Schuster et al., 2021), ASE (Moon et al., 2024), leaf, text width=45em ] ] [ Retrieval-Guided Initialization, fill=blue!10 [ RDM (Blattmann et al., 2022), kNN-Diffusion (Sheynin et al., 2022), Re-Imagen (Chen et al., 2022b),
ReDi (Zhang et al., 2023b), leaf, text width=45em ] ] ] ]
@@ 3.3.1 Efficient Solver
Given that the cost of sampling escalates proportionally with the number of discretized time steps, many researchers have concentrated on devising discretization schemes that reduce the number of time steps. A key insight emerges from reexamining the discrete forward process in the original DDPM formulation Eq.(1), as we reduce the step size between consecutive steps, the process naturally approaches a continuous transformation. Consequently, adopting learning-free methods using SDE or ODE solvers Song et al. (2020c) has been proposed.
SDE Solver. (Song et al., 2020c) firstly presents a stochastic differential equation (SDE) that smoothly transforms a complex data distribution to a known prior distribution by slowly injecting noise and a corresponding reverse-time SDE that transforms the prior distribution back into the data distribution by slowly removing the noise.The discrete noise addition steps in Eq.(1) are reformulated into a continuous process:
$d\mathbf{x}=\mathbf{f}(\mathbf{x},t)\,dt+g(t)\,d\bar{\mathbf{w}}$ (20)
$d\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\right]dt+g(t)d\bar{\mathbf{w}}$ (21)
Subsequently, there are many ways to efficiently implement SDE-based solvers. (Zhang & Chen, 2021)introduces a novel generative modeling and density estimation algorithm called Diffusion Normalizing Flow (DiffFlow). Similar to the SDE of diffusion models Eq.(20), the DiffFlow model also has a forward process:
$d\mathbf{x}=\mathbf{f}(\mathbf{x},t,\theta)dt+g(t)d\bar{\mathbf{w}}$ (22)
@@
$d\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t,\theta)-g^{2}(t)\mathbf{s}(\mathbf{x},t,\theta)\right]dt+g(t)d\bar{\mathbf{w}}$ (23)
As a result of the learnable parameter $\theta$, the drift term $f$ is also learnable in DiffFlow, compared to the fixed liner function as in most diffusion models. Besides, these SDEs are jointly trained by minimizing the KL divergence. This allows the model to better adapt to changes in the data distribution, thus speeding up the convergence of the backward diffusion process. Similar to DiffFlow,  Zand et al. (2023) proposes a method called Diffusion with Normalizing Flow priors that also combines diffusion models with normalizing flows. The method first uses a linear SDE in the forward process to gradually convert the data distribution into a noise distribution. In the reverse process, a normalizing flow network is introduced to map the standard Gaussian distribution to latent variables close to the data distribution through a series of reversible transformations, which allows the samples to return to the data distribution more quickly, rather than relying on a large number of small incremental adjustments.
@@
By optimizing the solving process of SDE in Eq.(20), Stochastic Adams Solver (SA-Solver) (Xue et al., 2024b) was presented. It is an innovative method designed to efficiently sample from Diffusion SDEs in Diffusion Probabilistic Models (DPMs) (Ho et al., 2020). By addressing the significant computational burden of traditional samplers, SA-Solver achieves this through a clever combination of variance-controlled diffusion SDEs and a stochastic Adams method (Buckwar & Winkler, 2006), which is a multi-step numerical technique that leverages prior evaluations to enhance efficiency. The method introduces a noise control function $\tau(t)$, enabling dynamic adjustment of the noise injected during sampling, which in turn strikes an optimal balance between sampling speed and the quality of the generated data. Operating within a predictor-corrector framework, SA-Solver first makes an initial estimate through the predictor step and then refines this estimate using the corrector step, ensuring greater accuracy in the final output. This strategic integration significantly reduces the number of function evaluations required.
ODE Solver. Unlike SDE solvers, the trajectories generated by ordinary differential equation (ODE) solvers are deterministic (Song et al., 2020c), remaining unaffected by stochastic variations. Consequently, these deterministic ODE solvers tend to achieve convergence more rapidly compared to their stochastic counterparts, although this often comes at the expense of a marginal reduction in sample quality. The corresponding deterministic process Eq.(24) can be derived from the reverse-time SDE Eq.(21) by removing the stochastic term $g(t)d\bar{\mathbf{w}}$, resulting in a deterministic process that shares the same marginal probability densities as the reverse-SDE:
$d\mathbf{x}=\left[{\mathbf{f}(\mathbf{x},t)}-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\right]dt$ (24)
$d\mathbf{x}=\mathbf{f}(\mathbf{x},t)dt$ (25)
Recent research has produced numerous works on faster diffusion samplers based on solving the ODE Eq.(24). Research shows that ODE samplers are highly effective when only a limited number of NFEs is available, while SDE samplers offer better resilience to prior mismatches (Nie et al., 2023) and exhibit superior performance with a greater number of NFEs (Lu et al., 2022).
Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020a) builds upon the framework of Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), offering significant enhancements in sampling efficiency, which is one of the first models to leverage ODEs explicitly for the accelerating sampling process.
@@
$d\mathbf{x}(t)=\epsilon_{\theta}^{(t)}\left(\frac{\mathbf{x}(t)}{\sqrt{\sigma^{2}+1}}\right)d\sigma(t)$ (28)
@@
To reduce the computational overhead,  Zheng et al. (2023) presents an improved technique for maximum likelihood estimation of ODEs. Instead of directly working with the drift and score terms in Eq.(24), it introduces velocity parameterization to predict and optimize velocity changes $d\mathbf{x}_{t}$ during the diffusion process directly. The method improves upon previous ODE-based approaches by incorporating second-order flow matching for more precise trajectory estimation. Additionally, it introduces a negative log-signal-to-noise-ratio (log-SNR) for timing control of the diffusion process, alongside normalized velocity and importance sampling to reduce variance and optimize training. These enhancements significantly improve the model’s likelihood estimation performance on image datasets without variational dequantization or data augmentation. While previous methods focus on improving reverse ODE integrators based on Eq.(24), Denoising MCMC (DMCMC) (Kim & Ye, 2022) takes a different approach by integrating Markov Chain Monte Carlo (MCMC) with ODE integrators to optimize the data sampling process. In DMCMC, MCMC first generates initialization points in the product space of data and diffusion time, which are closer to a noise-free state, significantly reducing the noise levels that need to be processed by the ODE integrators. This hybrid approach complements rather than improves the ODE integrators directly, enhancing overall sampling efficiency.
@@
$\frac{d\mathbf{x}_{t}}{dt}=\sigma_{d}\mathbf{F}_{\theta}(\frac{\mathbf{x}_{t}}{\sigma_{d}},c_{noise}(t))$ (29)
@@ 3.3.2 Sampling Scheduling
Parallel Sampling. Parallel sampling is a process that schedules sampling tasks in parallel. Traditional diffusion models require a extensive series of sequential denoising steps to generate a single sample, which can be quite slow. For instance, Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) might need thousands of these steps to produce one sample. However, parallel sampling leverages the power of a multi-core GPU to compute multiple sampling steps. This approach optimizes the use of computational resources and reduces the time needed for model generation. Currently, there is significant work on autoregressive models that employ parallelization to speed up the sampling process. However, these techniques cannot be directly applied to diffusion models. This is because the computational frameworks and inference efficiency in autoregressive models differ from those in diffusion models. Therefore, designing algorithms tailored to parallelize the sampling process of diffusion models is crucial.
@@
An innovative extension of the Denoising Diffusion Implicit Model (DDIM) (Song et al., 2020a) using Deep Equilibrium (DEQ) models is presented (Pokle et al., 2022), where the sampling sequence is conceptualized as a multivariate fixed-point system. This approach focuses on finding the system’s fixed point during the forward pass and utilizes implicit differentiation during the backward pass to enhance computational efficiency. By treating the sampling steps as an equilibrium system and solving for their fixed points simultaneously, parallel processing on multiple GPUs is achieved by batching the workload. Notably, it improves efficiency by updating each state $\mathbf{x}_{t}$ based on predictions from the noise prediction network $\epsilon_{\theta}$, which takes into account all subsequent states $\mathbf{x}_{t+1:T}$, unlike traditional diffusion processes that update states sequentially based only on the immediate next state $\mathbf{x}_{t+1}$.
@@
Building upon these parallel processing concepts, ParaTAA (Tang et al., 2024) also adopts an iterative approach, primarily applied in practical deployments for image generation tasks such as text-to-image transformations using Stable Diffusion. Specifically, ParaTAA enhances parallel sampling by solving triangular nonlinear equations through fixed-point iteration. Furthermore, the study introduces a novel variant of the Anderson Acceleration (Walker & Ni, 2011) technique, named Triangular Anderson Acceleration, designed to accelerate computation speed and improve the stability of iterative processes.
Timestep Schedule. In the sampling process of diffusion models, the entire process is discrete, and the model progressively restores data from noise through a series of discrete timesteps. Each timestep represents a small denoising step that moves the model from its current state closer to the real data. The timestep schedule refers to the strategy for selecting and arranging these timesteps.
@@
It may involve distributing them evenly or performing denser sampling during key stages to ensure the efficiency of the sampling process and the quality of the generated results. Selecting an appropriate method to choose a series of timesteps can enable the sampling process to converge quickly, which the process is shown in Figure 16.
@@
However, optimizing an exact Evidence Lower Bound (ELBO) is typically not conducive to enhancing image quality. To address this, a universal framework named Align Your Steps (AYS) (Sabour et al., 2024)has been introduced, aimed at optimizing sampling schedules for various datasets, models, and stochastic SDE solvers. AYS uses stochastic calculus to optimize the sampling schedule allowing for the identification of optimal sampling timesteps by minimizing the Kullback-Leibler Divergence Upper Bound (KLUB) between discretized learnt SDE and true learnt SDE, which is shown in Figure 17.

### v2

@@ 3.3 Efficient Sampling
The standard diffusion sampling process is computationally intensive because it requires sequentially executing a large number of denoising steps, with each step dependent on the output of the previous one as formulated in Eq(4). This inherent sequential dependency makes the sampling procedure time-consuming, resulting in significantly slower generation. To address these computational challenges, researchers have developed efficient sampling methods through four principal approaches. As illustrated in Figure 14, these encompass efficient SDE and ODE solvers that reduce the required number of function evaluations, advanced sampling scheduling strategies including parallel sampling techniques and timestep optimization methods, knowledge distillation techniques that transfer diffusion model capabilities to more efficient representations, and truncated sampling approaches that leverage early exit mechanisms and retrieval-based techniques to further accelerate the generation process while preserving output quality.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=12em,font=,, where level=2text width=16em,font=,, where level=3text width=24em,font=,, [ Efficient Sampling, ver [ Efficient Solver, fill=blue!10 [ SDE Solver, fill=blue!10 [ DiffFlow (Zhang & Chen, 2021), DiNoF (Zand et al., 2023),  Jolicoeur-Martineau et al. (2021),
GMS (Guo et al., 2024), BFM-Solver2 (Xue et al., 2024a), SA-Solver (Xue et al., 2024b),
Restart (Xu et al., 2023b), leaf, text width=45em ] ] [ ODE Solver, fill=blue!10 [ DDIM (Song et al., 2020a), DPM-solver (Lu et al., 2022), DEIS (Zhang & Chen, 2022),
@@
TAA (Walker & Ni, 2011), StreamDiffusion (Kodaira et al., 2023), leaf, text width=45em ] ] [ Timestep Scheduling, fill=blue!10 [ FastDPM (Kong & Ping, 2021),  Watson et al. (2021), AYS (Sabour et al., 2024), leaf, text width=45em ] ] ] [ Truncated Sampling, fill=blue!10 [ Early Exit, fill=blue!10 [ CATs (Schuster et al., 2021), ASE (Moon et al., 2024b), leaf, text width=45em ] ] [ Retrieval-Guided Initialization, fill=blue!10 [ RDM (Blattmann et al., 2022), kNN-Diffusion (Sheynin et al., 2022), Re-Imagen (Chen et al., 2022b),
ReDi (Zhang et al., 2023c), leaf, text width=45em ] ] ] [ Knowledge Distillation, fill=blue!10 [ Vector Field Distillation, fill=blue!10 [ Denoising Student (Luhman & Luhman, 2021), Progressive Distillation (Salimans & Ho, 2022),
 Meng et al. (2023), SANA-Sprint (Chen et al., 2025), DMD Yin et al. (2024), leaf, text width=45em ] ] [ Generator Distillation, fill=blue!10 [ SDS (Poole et al., 2022), VSD (Wang et al., 2024b), Diff-Instruct (Luo et al., 2024),
CSD (Decatur et al., 2024), leaf, text width=45em ] ] ] ]
@@ 3.3.1 Efficient Solver
As we mentioned in Section. §2.1.3, although diffusion solvers avoid the need to compute the intractable partition function, they still require numerous function evaluations with fixed step sizes for SDEs in Eq(11) or generic numerical integration methods for ODEs in Eq.(12), resulting in slow sampling procedures and substantial computational costs. These limitations create substantial computational costs and lengthy generation times. In the following sections, we present advances in both SDE and ODE methods that address these efficiency constraints.
SDE Solver. Nowadays, there are many ways to efficiently implement SDE-based solvers. (Zhang & Chen, 2021)introduces a novel generative modeling and density estimation algorithm called Diffusion Normalizing Flow (DiffFlow). Similar to the SDE of diffusion models Eq.(10), the DiffFlow model also has a forward process:
$\mathrm{d}\mathbf{x}=\mathbf{f}(\mathbf{x},t,\theta)\mathrm{d}t+g(t)\mathrm{d}\bar{\mathbf{w}}$ (32)
@@
$\mathrm{d}\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t,\theta)-g^{2}(t)\mathbf{s}(\mathbf{x},t,\theta)\right]\mathrm{d}t+g(t)\mathrm{d}\bar{\mathbf{w}}$ (33)
As a result of the learnable parameter $\theta$, the drift term $f$ is also learnable in DiffFlow, compared to the fixed liner function as in most diffusion models. Besides, these SDEs are jointly trained by minimizing the KL divergence. This allows the model to better adapt to changes in the data distribution, thus speeding up the convergence of the backward diffusion process. Similar to DiffFlow,  Zand et al. (2023) proposes a method called Diffusion with Normalizing Flow priors that also combines diffusion models with normalizing flows. The method first uses a linear SDE in the forward process to convert the data distribution into a noise distribution gradually. In the reverse process, a normalizing flow network is introduced to map the standard Gaussian distribution to latent variables close to the data distribution through a series of reversible transformations, which allows the samples to return to the data distribution more quickly, rather than relying on a large number of small incremental adjustments.
@@
By optimizing the solving process of SDE in Eq.(10), Stochastic Adams Solver (SA-Solver) (Xue et al., 2024b) was presented. It is an innovative method designed to efficiently sample from Diffusion SDEs in Diffusion Probabilistic Models (DPMs) (Ho et al., 2020). By addressing the significant computational burden of traditional samplers, SA-Solver achieves this through a clever combination of variance-controlled diffusion SDEs and a stochastic Adams method (Buckwar & Winkler, 2006), which is a multi-step numerical technique that leverages prior evaluations to enhance efficiency. The method introduces a noise control function $\tau(t)$, enabling dynamic adjustment of the noise injected during sampling, which in turn strikes an optimal balance between sampling speed and the quality of the generated data. Operating within a predictor-corrector framework, SA-Solver first makes an initial estimate through the predictor step and then refines this estimate using the corrector step, ensuring greater accuracy in the final output. This strategic integration significantly reduces the number of function evaluations required.  Xu et al. (2023b) introduces Restart, a novel sampling algorithm for diffusion models that strategically combines the strengths of SDE and ODE solvers. By theoretically analyzing Wasserstein bounds, the authors demonstrate that SDEs contract accumulated errors via stochasticity, while ODEs excel in low-NFE regimes due to smaller discretization errors. Restart bridges these regimes by alternating between noise injection and deterministic ODE backward steps, decoupling stochasticity from drift updates to amplify error contraction while retaining ODE efficiency.
ODE Solver. For efficiently implement the ODE-based solvers, Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020a) builds upon the framework of Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), offering significant enhancements in sampling efficiency, which is one of the first models to leverage ODEs explicitly for the accelerating sampling process.
@@
$\mathrm{d}\mathbf{x}(t)=\epsilon_{\theta}^{(t)}\left(\frac{\mathbf{x}(t)}{\sqrt{\sigma^{2}+1}}\right)\mathrm{d}\sigma(t)$ (36)
@@
To reduce the computational overhead,  Zheng et al. (2023) presents an improved technique for maximum likelihood estimation of ODEs. Instead of directly working with the drift and score terms in Eq.(12), it introduces velocity parameterization to predict and optimize velocity changes $\mathrm{d}\mathbf{x}_{t}$ during the diffusion process directly. The method improves upon previous ODE-based approaches by incorporating second-order flow matching for more precise trajectory estimation. Additionally, it introduces a negative log-signal-to-noise-ratio (log-SNR) for timing control of the diffusion process, alongside normalized velocity and importance sampling to reduce variance and optimize training. These enhancements significantly improve the model’s likelihood estimation performance on image datasets without variational dequantization or data augmentation. While previous methods focus on improving reverse ODE integrators based on Eq.(12), Denoising MCMC (DMCMC) (Kim & Ye, 2022) takes a different approach by integrating Markov Chain Monte Carlo (MCMC) with ODE integrators to optimize the data sampling process. In DMCMC, MCMC first generates initialization points in the product space of data and diffusion time, which are closer to a noise-free state, significantly reducing the noise levels that need to be processed by the ODE integrators. This hybrid approach complements rather than improves the ODE integrators directly, enhancing overall sampling efficiency.
@@
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\sigma_{d}\mathbf{F}_{\theta}(\frac{\mathbf{x}_{t}}{\sigma_{d}},c_{noise}(t))$ (37)
@@
Moreover,  (Du et al., 2024) replaces SDE with ODE’s deterministic trajectory, combined with fixed noise and Consistency Distillation Sampling loss, which minimizes stochastic fluctuations and computational redundancy. These designs enable faster convergence while maintaining high fidelity.
In conclusion, recent research has produced numerous works on faster diffusion samplers based on solving the ODE Eq.(12). Research shows that ODE samplers are highly effective when only a limited number of NFEs is available, while SDE samplers offer better resilience to prior mismatches (Nie et al., 2023) and exhibit superior performance with a greater Number of Function Evaluations (NFEs) (Lu et al., 2022).
@@ 3.3.2 Sampling Scheduling
Parallel Sampling. Parallel sampling is a process that schedules sampling tasks in parallel. Traditional diffusion models require a extensive series of sequential denoising steps to generate a single sample, which can be quite slow. For instance, Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) might need thousands of these steps to produce one sample. However, parallel sampling leverages the power of a multi-core GPU to compute multiple sampling steps. This approach optimizes the use of computational resources and reduces the time needed for model generation. Currently, there is significant work on autoregressive models that employ parallelization to speed up the sampling process.
@@
However, these techniques cannot be directly applied to diffusion models. This is because the computational frameworks and inference efficiency in autoregressive models differ from those in diffusion models. Therefore, designing algorithms tailored to parallelize the sampling process of diffusion models is crucial. An innovative extension of the Denoising Diffusion Implicit Model (DDIM) (Song et al., 2020a) using Deep Equilibrium (DEQ) models is presented (Pokle et al., 2022), where the sampling sequence is conceptualized as a multivariate fixed-point system. This approach focuses on finding the system’s fixed point during the forward pass and utilizes implicit differentiation during the backward pass to enhance computational efficiency. By treating the sampling steps as an equilibrium system and solving for their fixed points simultaneously, parallel processing on multiple GPUs is achieved by batching the workload. Notably, it improves efficiency by updating each state $\mathbf{x}_{t}$ based on predictions from the noise prediction network $\epsilon_{\theta}$, which takes into account all subsequent states $\mathbf{x}_{t+1:T}$, unlike traditional diffusion processes that update states sequentially based only on the immediate next state $\mathbf{x}_{t+1}$.
@@
Building upon these parallel processing concepts, ParaTAA (Tang et al., 2024) also adopts an iterative approach, primarily applied in practical deployments for image generation tasks such as text-to-image transformations using Stable Diffusion. Specifically, ParaTAA enhances parallel sampling by solving triangular nonlinear equations through fixed-point iteration. Furthermore, the study introduces a novel variant of the Anderson Acceleration (Walker & Ni, 2011) technique, named Triangular Anderson Acceleration, designed to accelerate computation speed and improve the stability of iterative processes.  Kodaira et al. (2023) introduces StreamDiffusion for real-time interactive diffusion by combining batched denoising to exploit GPU parallelism and Residual Classifier-Free Guidance(RCFG) to minimize redundant guidance computations, and input-output queues for asynchronous processing. It further optimizes energy efficiency via stochastic similarity filtering, which dynamically skips processing for near-identical frames using cosine similarity thresholds.
Timestep Schedule. In the sampling process of diffusion models, the entire process is discrete, and the model progressively restores data from noise through a series of discrete timesteps. Each timestep represents a small denoising step that moves the model from its current state closer to the real data. The timestep schedule refers to the strategy for selecting and arranging these timesteps. It may involve distributing them evenly or performing denser sampling during key stages to ensure the efficiency of the sampling process and the quality of the generated results. Selecting an appropriate method to choose a series of timesteps can enable the sampling process to converge quickly, which the process is shown in Figure 16.
@@
However, optimizing an exact Evidence Lower Bound (ELBO) is typically not conducive to enhancing image quality. To address this, Align Your Steps (AYS) (Sabour et al., 2024) proposes a compute-efficient paradigm for diffusion models by co-optimizing sampling schedules and solvers. Through stochastic calculus-guided Kullback-Leibler Divergence Upper Bound (KLUB) minimization, as shown in  17, AYS derives adaptive schedules that reduce sampling steps by 40% while maintaining quality—achieving equivalent FID scores with 20 steps versus 30 steps in baseline methods. The optimized schedules are versatile across modalities and solver types, enabling plug-and-play acceleration without model retraining.